how many 176 ohm resistors(in parallel) are required to carry 5A on a 220V line?

Grade:10

2 Answers

Aarti Gupta
8 years ago
Given
Current,I = 5A
Potential difference,V = 220V
Resistance of each resistors = 176ohm
No. of resistors connected in parallel = ?
Solution
Let total no. of resistors connected in parallel be ‘n’ so that total effective resistance be R.
We know that according to Ohm’s Law,
V = IR
Therefore, R = V/I = 220/5 = 44ohm
Now since n such resistors are connected in parallel,so
1/R = n/176
or, 1/44 = n/176
n = 176/44 = 4resistors
Thus 4 such resistors will be required to connected in parallel.
Adarsh
8 years ago
please check this answer